# Representing in Low Rank I: conjugacy, topological and homological aspects

### v1

@@ 1.3. Applications to the unit group of a group ring
The second application concerns the congruence kernel $C_{G}$ of the unit group of an integral group ring ${\mathcal{U}}({\mathbb{Z}}G)$. The latter has been investigated in [CdR1, CdR2, CdRZ]. In [CdR1, CdR2] several families of finite groups $G$ are described with the property that if $C_{G}$ is infinite, then $G$ belongs to one of these families. Under the assumption that ${\mathbb{Q}}G$ has no exceptional division components, an upper-bound on the virtual cohomological dimension of $C_{G}$ is obtained in [CdRZ, Theorem 5.1]. We contribute towards bounding from below, by showing that whenever ${\mathbb{Q}}G$ has an exceptional matrix component, then the congruence kernel contains a profinite free group of countable rank $\widehat{F_{\omega}}$. This is a direct consequence of [Lubot82, Theorem B], Theorem˜5.7 and the well-behavedness of the congruence kernel with direct products.
@@ 1.4. The blockwise subgroup isomorphism and Zassenhaus property
Recall that ${\mathcal{U}}({\mathbb{Z}}G)=\pm 1\cdot V({\mathbb{Z}}G)$ where $V({\mathbb{Z}}G)$ is the set of units in ${\mathcal{U}}({\mathbb{Z}}G)$ of augmentation $1$, called the normalised unit group. Three conjectures concerning finite subgroups $H$ of $V({\mathbb{Z}}G)$ were attributed to Zassenhaus [Zass, SehAntwerp]. Namely, consider the following sets:
@@
The $i$th Zassenhaus conjecture predicts that if $H\in\mathcal{G}_{i}$, then $H$ is ${\mathbb{Q}}G$-conjugated to a subgroup $K_{H}$ of $G$. Each of these three conjectures have been disproven [Scott, Rogg, EisMar], transforming them into the question for which $H$ and $G$ the above conjugation property holds. The above conjectures, but where conjugation over ${\mathbb{Q}}G$ is weakened to the existence of an abstract isomorphism between $H$ and some $K_{H}\leqslant G$, are called the subgroup isomorphism problem, see [Margolis, PdRV] and references therein.
The aim of Section˜7 is to study a blockwise variant of the Zassenhaus conjectures, i.e. investigating whether the image of $H$ under any irreducible $\mathbb{Q}$}-representation $\rho$ of $G$ is $\rho({\mathbb{Q}}G)$-conjugated to a subgroup of $\rho(G)$. Similarly, we also consider a blockwise subgroup isomorphism problem. Note that these problems can be investigated more generally. Therefore in Section˜7.1 we define the Zassenhaus and subgroup isomorphism property for general finite-dimensional semisimple $F$-algebras. The subgroup isomorphism property for the set $\mathcal{G}_{2}$ and twisted group rings has recently also received considerable attention, see [MaSc, MaSc2, MaSc3, HKSa, HKSi].
@@
We expect that combining the results from Section˜7.4 and the methods from [PdRV], one could obtain ˜G for any exceptional algebra.
@@ 1.5. The Virtual structure problem and rational isomorphism problem for exceptional components
In the sequel to this paper, [CJT], we will investigate the virtual structure problem and rational isomorphism problem for group algebras $FG$ having $\operatorname{(M_{exc})}$.
Recall that in [KleinertSurvey], Kleinert formulated formally the problem to obtain unit theorems for the unit group of orders $\mathcal{O}$ in finite-dimensional semisimple algebras $A$. More precisely, the problem concerns finding classes of groups $\mathcal{G}(A)$ such that most of the generic unit groups of $A$ are contained in $\mathcal{G}(A)$. Formulated a bit differently, one could consider a class of groups $\mathcal{G}$ and classify all semisimple algebras $A$ such that $A$ has a unit theorem for $\mathcal{G}$. Now if we restrict ourselves to group algebras, then the simple factors are related to each other via the representation theory of the group basis. As such the following problem was distilled in [JesRioReine].
@@
For a historical overview on the virtual structure problem we refer the reader to the introduction of [J1].
A first main result resolves the virtual structure problem for the class of $\operatorname{(M_{exc})}$ groups. This result completes a line of research started $30$ years ago with among others the papers [JesLealRio, JesRioReine, PitaRioRuiz, FreebyFree]. The most general result until now concerned the finite groups $G$ such that ${\mathbb{Q}}G$ has the stronger property that
@@
which were classified in [JesRioCrelle, FreebyFree]. Considering property $\operatorname{(M_{exc})}$ generalises their considerations in two ways. Anecdotally, the MathSciNet review of [JesRioCrelle] states that they obtain “the most far-reaching result possible”, which indeed looked like to be so for now almost 20 years. We will prove the following.
@@
Finally, we study the rational isomorphism problem, i.e. the problem of distinguishing a finite group among all finite groups through its rational group algebra. Positive instances of the latter problem are scarce as the authors are only aware of the case of abelian groups [PW] and metacyclic groups [GBdR1, GBdR2]. Note that Dade’s examples [Dade] are metabelian, indicating the difficulty to find new classes. Although finite groups satisfying $\operatorname{(M_{exc})}$ are metabelian by ˜H, we will show that they are distinguishable.
@@ 2. The block Virtual Structure Problem
Note that $\operatorname{RNr}_{A/K}(\cdot)$ is a multiplicative map, $\operatorname{RNr}_{A/K}(A)\subseteq K$ and $\operatorname{RNr}_{A/K}(a)$ only depends on $K$ and $a\in A$ (and not on the chosen splitting field $E$ or the isomorphism $E\otimes_{K}A\cong\operatorname{M}_{n}(E)$), see [EricAngel1, page 51]. For a subring $R$ of $A$, define
@@
for the class of groups which are formed out of direct products of groups in $\mathcal{G}$ and abelian groups. Using [KleRio], together with classical results on unit groups of orders, we obtain the following. We will say that a group is virtually-$\mathcal{G}$ if it has a finite index subgroup which belongs to the class $\mathcal{G}$.
@@
It was proven by Kleinert that $\operatorname{SL}_{1}(\mathcal{O})$ (for $\mathcal{O}$ an order in a finite-dimensional simple $\mathbb{Q}$}-algebra AA) is finite if and only if $A$ is a field or a totally definite quaternion algebra, see [Kleinert] or [EricAngel1, Proposition 5.5.6]. Moreover, a classification of the rings of integers $R$ and finite groups $G$ such that ${\mathcal{U}}(RG)$ is virtually cyclic was obtained in [J1, Proposition 1.4].
@@
The proof of Lemma˜2.1 will show that the same conclusion holds if the class of groups that are virtually-$\prod\mathcal{G}$ has the following weaker direct factor property: if $\prod_{e\in\operatorname{PCI}(A)}\operatorname{SL}_{1}(\mathcal{O}e)\times{\mathcal{U}}(\mathcal{Z}(\mathcal{O}))$ is virtually-$\prod\mathcal{G}$, then each $\operatorname{SL}_{1}(\mathcal{O}e)$ is virtually-$\prod\mathcal{G}$. Or more abstractly, thanks to [Kleinert], it is enough to obtain the property for direct products of the form $\prod_{i=1}^{m}\Gamma_{i}\times\Gamma$ with $\Gamma_{i}$ non-abelian virtually indecomposable and $\Gamma$ a finitely generated abelian group.
@@
Let $\mathcal{G}$ be the class consisting of groups $\Gamma$ having a (potentially trivial) normal subgroup $N$ which is free such that $\Gamma/N$ is also free. Then it was proven in [FreebyFree, Theorem 2.1] that $\mathcal{G}$ satisfies the hypothesis from Lemma˜2.1.
Another example is the class of groups having infinitely many ends. By Stalling’s theorem this is equivalent to having an amalgamated or HNN splitting over a finite group. In [J1, Lemma 2.1] it is proven that this class also satisfies the hypothesis from Lemma˜2.1.
@@
We know that ${\mathcal{U}}(\mathcal{O}_{1})$ and ${\mathcal{U}}(\mathcal{O}_{2})$ are commensurable. Therefore [EricAngel1, Proposition 5.5.1] implies that the subgroups $\langle\operatorname{SL}_{1}(\mathcal{O}_{j}),{\mathcal{U}}(\mathcal{Z}(\mathcal{O}_{j}))\rangle$ for $j=1,2$ are commensurable. Furthermore, [EricAngel1, Corollary 5.5.3] implies that $\langle\operatorname{SL}_{1}(\mathcal{O}_{j}),{\mathcal{U}}(\mathcal{Z}(\mathcal{O}_{j}))\rangle$ is commensurable with $\operatorname{SL}_{1}(\mathcal{O}_{j})\times{\mathcal{U}}(\mathcal{Z}(\mathcal{O}_{j}))$. From this we conclude that the $\operatorname{SL}_{1}(\mathcal{O}_{j})$, $j=1,2$ are necessarily commensurable. ∎
@@
The sufficiency in (1) is clear. Hence assume that ${\mathcal{U}}(\mathcal{O})$ is virtually-$\prod\mathcal{G}$. Recall [EricAngel1, Proposition 5.5.1], which says that $\langle\operatorname{SL}_{1}(\mathcal{O}),{\mathcal{U}}(\mathcal{Z}(\mathcal{O}))\rangle$ is a subgroup of finite index in ${\mathcal{U}}(\mathcal{O})$ and moreover that $\operatorname{SL}_{1}(\mathcal{O})\cap{\mathcal{U}}(\mathcal{Z}(\mathcal{O}))$ is finite. Therefore, $\operatorname{SL}_{1}(\mathcal{O})\times{\mathcal{U}}(\mathcal{Z}(\mathcal{O}))$ is commensurable with ${\mathcal{U}}(\mathcal{O})$. Furthermore, $\operatorname{SL}_{1}(\mathcal{O})$ is commensurable with $\prod_{e\in\operatorname{PCI}(A)}\operatorname{SL}_{1}(\mathcal{O}e)$ (see also [EricAngel1, Corollary 5.5.3]). Hence by the assumptions on $\mathcal{G}$, we have that $\prod_{e\in\operatorname{PCI}(A)}\operatorname{SL}_{1}(\mathcal{O}e)\times{\mathcal{U}}(\mathcal{Z}(\mathcal{O}))$ and subsequently $\operatorname{SL}_{1}(\mathcal{O}e)$ is virtually-$\prod\mathcal{G}$ for each $e\in\operatorname{PCI}(A)$. Now if $Ae$ is a field or a totally definite quaternion algebra, then $\operatorname{SL}_{1}(\mathcal{O}e)$ is finite by [Kleinert]. For other simple algebras $Ae$, [KleRio, Theorem 1] says that $\operatorname{SL}_{1}(\mathcal{O}e)$ is virtually indecomposable, meaning that if $\operatorname{SL}_{1}(\mathcal{O}e)$ is commensurable with a direct product $\Gamma_{1}\times\Gamma_{2}$, then $\Gamma_{1}$ or $\Gamma_{2}$ is finite. Therefore, $\operatorname{SL}_{1}(\mathcal{O}e)$ being virtually-$\prod\mathcal{G}$ implies that it is either virtually cyclic or virtually-$\mathcal{G}$, as desired.
For part (2) we may assume without loss of generality that $A$ is simple. It follows from [JanJesTem, Theorem 4.1] that ${\mathcal{U}}(\mathcal{O})$ contains a free group if and only if $A$ is not a totally definite quaternion algebra. In particular, combined with [EricAngel1, Proposition 5.5.6], (2) follows.
Now we consider (3). By the first part, $\operatorname{SL}_{1}(\mathcal{O}e)$ is either virtually cyclic or virtually-$\mathcal{G}$. Part (2) and [EricAngel1, Proposition 5.5.6] moreover imply that the former case only happens if $Ae$ is a field or a totally definite quaternion algebra. In particular the degree of $Ae$ is at most $2$. Hence suppose that $\operatorname{SL}_{1}(\mathcal{O}e)$ is virtually-$\mathcal{G}$. In that case, the desired conclusion follows from [KleRio, Theorem 3]. ∎
@@ 3.1. The role of the coefficient field
This shows the importance of first understanding the case of rational group algebras. In the companion paper [CJT], where the finite groups such that $FG$ has $\operatorname{(M_{exc})}$ will be classified, the bulk of the work will be dedicated to the case $F={\mathbb{Q}}$.
@@
Finite groups $G$ for which ${\mathbb{Q}}G$ has only division algebra components have been classified in [SehNilpotent, Theorem 3.5]. Concretely, ${\mathbb{Q}}G$ is a direct product of division algebras if and only if $G$ is abelian or $G$ is of the form $A\times C_{2}^{n}\times Q_{8}$ with $A$ an abelian group of odd order such that the multiplicative order of $2\in{\mathbb{Q}}$ is odd modulo $|A|$.
@@
Recall that there is a bijection between the simple components of $LG$ and the absolutely irreducible characters $\operatorname{Irr}(G)$ of $G$. For $\chi\in\operatorname{Irr}(G)$ denote the associated component of $LG$ by $A_{L}(\chi)$. Now, following [EricAngel1, Theorem 3.3.1], there is an isomorphism of $F(\chi)$-algebras

### v2

@@ 1.3. Applications to the unit group of a group ring
The second application concerns the congruence kernel $C_{G}$ of the unit group of an integral group ring ${\mathcal{U}}({\mathbb{Z}}G)$. The latter has been investigated in [9, 10, 13]. In [9, 10] several families of finite groups $G$ are described with the property that if $C_{G}$ is infinite, then $G$ belongs to one of these families. Under the assumption that ${\mathbb{Q}}G$ has no exceptional division components, an upper-bound on the virtual cohomological dimension of $C_{G}$ is obtained in [13, Theorem 5.1]. We contribute towards bounding from below, by showing that whenever ${\mathbb{Q}}G$ has an exceptional matrix component, then the congruence kernel contains a profinite free group of countable rank $\widehat{F_{\omega}}$. This is a direct consequence of [58, Theorem B], Theorem˜5.7 and the well-behavedness of the congruence kernel with direct products.
@@ 1.4. The blockwise subgroup isomorphism and Zassenhaus property
Recall that ${\mathcal{U}}({\mathbb{Z}}G)=\pm 1\cdot V({\mathbb{Z}}G)$ where $V({\mathbb{Z}}G)$ is the set of units in ${\mathcal{U}}({\mathbb{Z}}G)$ of augmentation $1$, called the normalised unit group. Three conjectures concerning finite subgroups $H$ of $V({\mathbb{Z}}G)$ were attributed to Zassenhaus [88, 80]. Namely, consider the following sets:
@@
The $i$th Zassenhaus conjecture predicts that if $H\in\mathcal{G}_{i}$, then $H$ is ${\mathbb{Q}}G$-conjugated to a subgroup $K_{H}$ of $G$. Each of these three conjectures have been disproven [77, 74, 24], transforming them into the question for which $H$ and $G$ the above conjugation property holds. The above conjectures, but where conjugation over ${\mathbb{Q}}G$ is weakened to the existence of an abstract isomorphism between $H$ and some $K_{H}\leqslant G$, are called the subgroup isomorphism problem, see [60, 68] and references therein.
The aim of Section˜7 is to study a blockwise variant of the Zassenhaus conjectures, i.e. investigating whether the image of $H$ under any irreducible $\mathbb{Q}$}-representation $\rho$ of $G$ is $\rho({\mathbb{Q}}G)$-conjugated to a subgroup of $\rho(G)$. Similarly, we also consider a blockwise subgroup isomorphism problem. Note that these problems can be investigated more generally. Therefore in Section˜7.1 we define the Zassenhaus and subgroup isomorphism property for general finite-dimensional semisimple $F$-algebras. The subgroup isomorphism property for the set $\mathcal{G}_{2}$ and twisted group rings has recently also received considerable attention, see [63, 64, 65, 35, 51].
@@
We expect that combining the results from Section˜7.4 and the methods from [68], one could obtain ˜G for any exceptional algebra.
@@ 1.5. The Virtual structure problem and rational isomorphism problem for exceptional components
In the sequel to this paper, [15], we will investigate the virtual structure problem and rational isomorphism problem for group algebras $FG$ having $\operatorname{(M_{exc})}$.
Recall that in [56], Kleinert formulated formally the problem to obtain unit theorems for the unit group of orders $\mathcal{O}$ in finite-dimensional semisimple algebras $A$. More precisely, the problem concerns finding classes of groups $\mathcal{G}(A)$ such that most of the generic unit groups of $A$ are contained in $\mathcal{G}(A)$. Formulated a bit differently, one could consider a class of groups $\mathcal{G}$ and classify all semisimple algebras $A$ such that $A$ has a unit theorem for $\mathcal{G}$. Now if we restrict ourselves to group algebras, then the simple factors are related to each other via the representation theory of the group basis. As such the following problem was distilled in [46].
@@
For a historical overview on the virtual structure problem we refer the reader to the introduction of [42].
A first main result resolves the virtual structure problem for the class of $\operatorname{(M_{exc})}$ groups. This result completes a line of research started $30$ years ago with among others the papers [48, 46, 71, 50]. The most general result until now concerned the finite groups $G$ such that ${\mathbb{Q}}G$ has the stronger property that
@@
which were classified in [46, 50]. Considering property $\operatorname{(M_{exc})}$ generalises their considerations in two ways. Anecdotally, the MathSciNet review of [46] states that they obtain “the most far-reaching result possible”, which indeed looked like to be so for now almost 20 years. We will prove the following.
@@
Finally, we study the rational isomorphism problem, i.e. the problem of distinguishing a finite group among all finite groups through its rational group algebra. Positive instances of the latter problem are scarce as the authors are only aware of the case of abelian groups [70] and metacyclic groups [29, 30]. Note that Dade’s examples [16] are metabelian, indicating the difficulty to find new classes. Although finite groups satisfying $\operatorname{(M_{exc})}$ are metabelian by ˜H, we will show that they are distinguishable.
@@ 2. The block Virtual Structure Problem
Note that $\operatorname{RNr}_{A/K}(\cdot)$ is a multiplicative map, $\operatorname{RNr}_{A/K}(A)\subseteq K$ and $\operatorname{RNr}_{A/K}(a)$ only depends on $K$ and $a\in A$ (and not on the chosen splitting field $E$ or the isomorphism $E\otimes_{K}A\cong\operatorname{M}_{n}(E)$), see [47, page 51]. For a subring $R$ of $A$, define
@@
for the class of groups which are formed out of direct products of groups in $\mathcal{G}$ and abelian groups. Using [55], together with classical results on unit groups of orders, we obtain the following. We will say that a group is virtually-$\mathcal{G}$ if it has a finite index subgroup which belongs to the class $\mathcal{G}$.
@@
It was proven by Kleinert that $\operatorname{SL}_{1}(\mathcal{O})$ (for $\mathcal{O}$ an order in a finite-dimensional simple $\mathbb{Q}$}-algebra AA) is finite if and only if $A$ is a field or a totally definite quaternion algebra, see [54] or [47, Proposition 5.5.6]. Moreover, a classification of the rings of integers $R$ and finite groups $G$ such that ${\mathcal{U}}(RG)$ is virtually cyclic was obtained in [42, Proposition 1.4].
@@
The proof of Lemma˜2.1 will show that the same conclusion holds if the class of groups that are virtually-$\prod\mathcal{G}$ has the following weaker direct factor property: if $\prod_{e\in\operatorname{PCI}(A)}\operatorname{SL}_{1}(\mathcal{O}e)\times{\mathcal{U}}(\mathcal{Z}(\mathcal{O}))$ is virtually-$\prod\mathcal{G}$, then each $\operatorname{SL}_{1}(\mathcal{O}e)$ is virtually-$\prod\mathcal{G}$. Or more abstractly, thanks to [54], it is enough to obtain the property for direct products of the form $\prod_{i=1}^{m}\Gamma_{i}\times\Gamma$ with $\Gamma_{i}$ non-abelian virtually indecomposable and $\Gamma$ a finitely generated abelian group.
@@
Let $\mathcal{G}$ be the class consisting of groups $\Gamma$ having a (potentially trivial) normal subgroup $N$ which is free such that $\Gamma/N$ is also free. Then it was proven in [50, Theorem 2.1] that $\mathcal{G}$ satisfies the hypothesis from Lemma˜2.1.
Another example is the class of groups having infinitely many ends. By Stalling’s theorem this is equivalent to having an amalgamated or HNN splitting over a finite group. In [42, Lemma 2.1] it is proven that this class also satisfies the hypothesis from Lemma˜2.1.
@@
We know that ${\mathcal{U}}(\mathcal{O}_{1})$ and ${\mathcal{U}}(\mathcal{O}_{2})$ are commensurable. Therefore [47, Proposition 5.5.1] implies that the subgroups $\langle\operatorname{SL}_{1}(\mathcal{O}_{j}),{\mathcal{U}}(\mathcal{Z}(\mathcal{O}_{j}))\rangle$ for $j=1,2$ are commensurable. Furthermore, [47, Corollary 5.5.3] implies that $\langle\operatorname{SL}_{1}(\mathcal{O}_{j}),{\mathcal{U}}(\mathcal{Z}(\mathcal{O}_{j}))\rangle$ is commensurable with $\operatorname{SL}_{1}(\mathcal{O}_{j})\times{\mathcal{U}}(\mathcal{Z}(\mathcal{O}_{j}))$. From this we conclude that the $\operatorname{SL}_{1}(\mathcal{O}_{j})$, $j=1,2$ are necessarily commensurable. ∎
@@
The sufficiency in (1) is clear. Hence assume that ${\mathcal{U}}(\mathcal{O})$ is virtually-$\prod\mathcal{G}$. Recall [47, Proposition 5.5.1], which says that $\langle\operatorname{SL}_{1}(\mathcal{O}),{\mathcal{U}}(\mathcal{Z}(\mathcal{O}))\rangle$ is a subgroup of finite index in ${\mathcal{U}}(\mathcal{O})$ and moreover that $\operatorname{SL}_{1}(\mathcal{O})\cap{\mathcal{U}}(\mathcal{Z}(\mathcal{O}))$ is finite. Therefore, $\operatorname{SL}_{1}(\mathcal{O})\times{\mathcal{U}}(\mathcal{Z}(\mathcal{O}))$ is commensurable with ${\mathcal{U}}(\mathcal{O})$. Furthermore, $\operatorname{SL}_{1}(\mathcal{O})$ is commensurable with $\prod_{e\in\operatorname{PCI}(A)}\operatorname{SL}_{1}(\mathcal{O}e)$ (see also [47, Corollary 5.5.3]). Hence by the assumptions on $\mathcal{G}$, we have that $\prod_{e\in\operatorname{PCI}(A)}\operatorname{SL}_{1}(\mathcal{O}e)\times{\mathcal{U}}(\mathcal{Z}(\mathcal{O}))$ and subsequently $\operatorname{SL}_{1}(\mathcal{O}e)$ is virtually-$\prod\mathcal{G}$ for each $e\in\operatorname{PCI}(A)$. Now if $Ae$ is a field or a totally definite quaternion algebra, then $\operatorname{SL}_{1}(\mathcal{O}e)$ is finite by [54]. For other simple algebras $Ae$, [55, Theorem 1] says that $\operatorname{SL}_{1}(\mathcal{O}e)$ is virtually indecomposable, meaning that if $\operatorname{SL}_{1}(\mathcal{O}e)$ is commensurable with a direct product $\Gamma_{1}\times\Gamma_{2}$, then $\Gamma_{1}$ or $\Gamma_{2}$ is finite. Therefore, $\operatorname{SL}_{1}(\mathcal{O}e)$ being virtually-$\prod\mathcal{G}$ implies that it is either virtually cyclic or virtually-$\mathcal{G}$, as desired.
For part (2) we may assume without loss of generality that $A$ is simple. It follows from [44, Theorem 4.1] that ${\mathcal{U}}(\mathcal{O})$ contains a free group if and only if $A$ is not a totally definite quaternion algebra. In particular, combined with [47, Proposition 5.5.6], (2) follows.
Now we consider (3). By the first part, $\operatorname{SL}_{1}(\mathcal{O}e)$ is either virtually cyclic or virtually-$\mathcal{G}$. Part (2) and [47, Proposition 5.5.6] moreover imply that the former case only happens if $Ae$ is a field or a totally definite quaternion algebra. In particular the degree of $Ae$ is at most $2$. Hence suppose that $\operatorname{SL}_{1}(\mathcal{O}e)$ is virtually-$\mathcal{G}$. In that case, the desired conclusion follows from [55, Theorem 3]. ∎
@@ 3.1. The role of the coefficient field
This shows the importance of first understanding the case of rational group algebras. In the companion paper [15], where the finite groups such that $FG$ has $\operatorname{(M_{exc})}$ will be classified, the bulk of the work will be dedicated to the case $F={\mathbb{Q}}$.
@@
Finite groups $G$ for which ${\mathbb{Q}}G$ has only division algebra components have been classified in [79, Theorem 3.5]. Concretely, ${\mathbb{Q}}G$ is a direct product of division algebras if and only if $G$ is abelian or $G$ is of the form $A\times C_{2}^{n}\times Q_{8}$ with $A$ an abelian group of odd order such that the multiplicative order of $2\in{\mathbb{Q}}$ is odd modulo $|A|$.
@@
Recall that there is a bijection between the simple components of $LG$ and the absolutely irreducible characters $\operatorname{Irr}(G)$ of $G$. For $\chi\in\operatorname{Irr}(G)$ denote the associated component of $LG$ by $A_{L}(\chi)$. Now, following [47, Theorem 3.3.1], there is an isomorphism of $F(\chi)$-algebras